# Mathematical Anatomy of Neutrino Decoherence in Red Turbulence: A Fractional Calculus Approach

### v1

@@ Abstract
We develop an exact framework for neutrino decoherence in power-law correlated turbulent matter, as encountered in core-collapse supernovae. Employing the Nakajima–Zwanzig projection technique, we derive an exact non-Markovian master equation for the neutrino density matrix. For red turbulence spectra, the memory kernel follows $K(t)\propto t^{-\nu}$, leading naturally to a fractional calculus description. To maintain mathematical rigor for $\nu\geq 1$, we introduce an exponential cutoff at the dissipation scale, regularizing ultraviolet divergences.
The exact Laplace-space solution for the survival probability is obtained. In the high-density matter basis relevant to supernovae, the solution is expressed through Mittag-Leffler functions, establishing a direct link to anomalous diffusion phenomena. For extremely red spectra ($\nu<0$), the memory integral corresponds to a higher-order fractional operator.
Our work clarifies how spectral index, regularization scale, and decoherence efficiency interrelate, providing a complete analytical description and practical tools for supernova neutrino simulations. The fractional calculus formulation reveals fundamental mathematical connections between neutrino flavor evolution and other systems governed by long-range temporal correlations.
@@ I Introduction
Core-collapse supernovae (CCSNe) represent one of the most energetic and complex astrophysical phenomena in the Universe, marking the cataclysmic death of massive stars ($M\gtrsim 8M_{\odot}$). These explosions not only seed the interstellar medium with heavy elements essential for planetary formation and life, but also give birth to neutron stars and black holes [28, 12, 39]. The detection of gravitational waves from binary neutron star mergers [2, 3] and the anticipated observation of neutrinos from the next Galactic supernova by next-generation detectors such as Hyper-Kamiokande [27] and DUNE [5] underscore the importance of understanding these events through multi-messenger astronomy.
@@
Despite decades of theoretical and computational advancements, the precise explosion mechanism of CCSNe remains an open question [29, 13, 24]. Current three-dimensional simulations suggest that the neutrino-heating mechanism, aided by hydrodynamic instabilities such as the standing accretion shock instability (SASI) and neutrino-driven convection, can successfully revive the stalled shock in many progenitor models [4, 16]. Turbulence plays a crucial role in this process by enhancing the efficiency of neutrino heating and facilitating shock revival [15, 14]. Recent high-resolution simulations reveal that turbulent energy cascades follow Kolmogorov-like scaling in the inertial range [34], but the precise characterization of turbulence in supernova environments remains challenging [40].
@@
Neutrinos, carrying away $\sim 99\%$ of the gravitational binding energy ($\sim 3\times 10^{53}$ erg), are the primary cooling agents and potential drivers of the explosion through the neutrino-heating mechanism [28, 38]. Their flavor evolution is profoundly affected by both coherent forward scattering (the Mikheyev-Smirnov-Wolfenstein (MSW) effect) [36, 45] and collective oscillations due to neutrino-neutrino interactions [20, 19, 26]. Recent studies have identified fast flavor instabilities that can occur on centimeter scales and nanosecond timescales [8, 17, 21, 23], potentially altering the neutrino energy deposition and nucleosynthesis yields [41, 22, 44]. Moreover, muon creation in supernova matter has been shown to facilitate neutrino-driven explosions [9], adding further complexity to the flavor evolution landscape.
@@
In addition to vacuum and matter effects, turbulence in the supernova envelope introduces stochastic density fluctuations that can significantly impact neutrino flavor evolution. The pioneering work of Loreti and Balantekin [35] established the framework for treating neutrino oscillations in noisy media, demonstrating that density fluctuations can induce decoherence and modify survival probabilities. Subsequent studies have extended this analysis to various turbulent environments [33, 43, 31, 42]. For power-law correlated turbulence, the master equation acquires a non-Markovian character with memory kernels scaling as $K(t)\propto t^{-\nu}$, where $\nu$ is the spectral index determined by the turbulence power spectrum [32, 10, 1]. Recently, Mukhopadhyay and Sen [37] investigated how turbulence signatures in supernova neutrinos could be probed in upcoming detectors, highlighting the observational relevance of this phenomenon.
@@
The mathematical treatment of neutrino decoherence in power-law correlated turbulent media presents significant challenges due to the non-Markovian nature of the evolution equations. The master equation with memory kernel $K(t)\propto t^{-\nu}$ naturally connects to fractional calculus, where the time convolution integral represents a fractional integral operator of order $1-\nu$ [30]. For $\nu<1$, this corresponds to the Riemann-Liouville fractional integral, while for $\nu\geq 1$, regularization at small time scales becomes necessary to ensure mathematical consistency. Fractional calculus has found widespread application in describing anomalous diffusion, viscoelastic materials, and non-Markovian processes in open quantum systems [11]. In the context of neutrino physics, the emergence of Mittag-Leffler functions $E_{\alpha,\beta}(z)$ as natural solutions to fractional differential equations provides a powerful analytical framework [18].
@@
While approximate and numerical treatments of neutrino propagation in turbulent media exist [35, 32]—ranging from white-noise models to power-law correlated fluctuations (numerical)—a complete analytical solution for the survival probability with generic power-law memory kernels has remained elusive. Previous approaches often relied on perturbative expansions or specialized limits, leaving the full mathematical structure unexplored. The power-law form of the memory kernel, $K(t)\propto t^{-\nu}$, suggests a natural connection to fractional calculus, yet this connection has not been fully exploited to obtain closed-form analytical solutions.
@@
Several critical gaps in the current understanding warrant systematic investigation: (i) the explicit formulation of the neutrino propagation problem within a fractional calculus framework; (ii) the mathematical regularization required for spectral exponents $\nu\geq 1$ to handle ultraviolet divergences at short time scales ($t\to 0$); (iii) the derivation of exact analytical solutions for arbitrary $\nu$, expressed through special functions like Mittag-Leffler functions; (iv) the distinctive mathematical features arising for negative spectral exponents ($\nu<0$) corresponding to extremely red spectra; (v) the transition between algebraic and exponential decoherence regimes controlled by regularization; and (vi) the establishment of explicit connections between neutrino decoherence and anomalous diffusion phenomena through their shared fractional calculus structure.
@@
This work addresses these gaps by developing a comprehensive exact theory based on fractional calculus. We establish the precise correspondence between the neutrino master equation with power-law memory and fractional differential equations. Special attention is paid to the ultraviolet divergence of the correlation function for $\nu\geq 1$, which we regularize by introducing an exponential cutoff $\exp(-\epsilon/t)$ at the dissipation time scale $\epsilon$, ensuring mathematical consistency while preserving the essential power-law behavior in the inertial range. This regularization enables the derivation of exact, closed-form solutions valid for all $\nu$.
@@
The exact solution reveals a rich mathematical structure: for $\nu<1$, the memory integral corresponds to a Riemann-Liouville fractional integral of order $1-\nu$, while for $\nu<0$, it becomes a higher-order fractional operator. The survival probability is expressed exactly in terms of Mittag-Leffler functions, which naturally interpolate between initial algebraic decay and eventual exponential relaxation. This formulation provides not only a complete analytical description but also efficient computational methods through established properties of fractional calculus and special functions.
@@
Our work demonstrates that the fractional calculus approach offers a unifying mathematical framework that connects neutrino decoherence in turbulent matter with broader phenomena in statistical physics, including anomalous diffusion and viscoelastic relaxation. This perspective yields new insights into the interplay between spectral index, regularization scale, and decoherence efficiency, providing both fundamental understanding and practical tools for supernova neutrino simulations.
@@
In this work, we establish the complete mathematical formalism for the exact theory of neutrino decoherence in red turbulence. We provide a rigorous derivation of the master equation for the neutrino density matrix via an open quantum systems approach, obtaining the exact non-Markovian equation through the Nakajima–Zwanzig projection technique [11]. Furthermore, we derive the exact solution for the electron neutrino survival probability in Laplace space. For red spectral turbulence, we develop a fractional calculus framework, yielding a fractional differential equation exactly solvable in terms of Mittag-Leffler functions. This fractional formulation not only provides efficient computational methods but also uncovers profound mathematical links between neutrino flavor evolution and other physical systems characterized by power-law correlations and anomalous dynamics.
@@
This paper is organized as follows: Section II reviews the exact Laplace-space solution with small-scale regularization. Section III develops the analytic solution via fractional calculus. Section IV summarizes our findings and outlines directions for future research. In a companion paper [7], we extend this analysis to include collective neutrino oscillations in turbulent environments.
@@ II.1 Physical Picture and Assumptions
We begin by outlining the physical scenario and the key assumptions underlying our derivation. The propagation of neutrinos through turbulent matter in core-collapse supernovae involves multiple scales:
  • Microscopic scale ($\sim 10^{-12}$ cm): Neutrino flavor oscillations governed by quantum mechanics.
  • Turbulence inertial range ($10^{2}$$10^{6}$ cm): Compressible hydrodynamic turbulence with power-law scaling [25].
  • Supernova scale ($\sim 10^{8}$ cm): Overall system size.
@@ II.2.1 Obukhov-Corrsin Spectrum and Dimensionless Formulation
Let us begin by introducing dimensionless variables for a formal simplification. Let $n_{0}$ be a reference electron number density (typically the mean density $\bar{n}_{e}$), and $T_{0}$ a reference time (to be chosen as $1/\Delta_{m}$, where $\Delta_{m}$ is the matter-level splitting of neutrinos). In natural units ($\hbar=c=1$), length, time, and energy satisfy $L]=[T]=[E]^{-1}$. We define dimensionless coordinates and density fluctuations:
@@
which is itself dimensionless. For fully developed power-law turbulence, the correlation function exhibits power-law scaling in the inertial range, but requires regularization at both small and large scales. While large-scale correlations are naturally limited by the system size, the small-scale behavior requires careful treatment due to potential divergences at $\tau\to 0$.
II.2.2 Small-Scale Regularized Correlation Function
The power-law correlation $K(\tau)\propto\tau^{-\nu}$ diverges as $\tau\to 0$ for $\nu\geq 1$, representing an unphysical ultraviolet divergence. In real turbulent systems, this divergence is regularized by physical processes at the dissipation scale. We introduce a small-scale cutoff at the dissipation time scale $\tau_{\text{diss}}=l_{\text{diss}}/c$, where $l_{\text{diss}}$ is the Kolmogorov dissipation scale:
@@ II.2.2 Small-Scale Regularized Correlation Function
For $\nu<1$, the correlation function is integrable at $\tau=0$ even without regularization, but we maintain the same form for consistency. The normalization constant $A$ is chosen such that $\int_{0}^{\infty}\tilde{\mathcal{C}}(\tilde{\tau})d\tilde{\tau}=1$ when considering the effective damping rate.
@@ II.2.3 Exact Non-Markovian Master Equation from Nakajima–Zwanzig Projection
We now present a rigorous, non-perturbative derivation of the master equation governing neutrino decoherence in red turbulent matter using the Nakajima–Zwanzig (NZ) projection operator formalism [11]. The essential advantage of the NZ approach is that it treats the memory kernel exactly for classical Gaussian noise, automatically resumming all cumulants that contribute to the phase‑diffusion process.
@@
$-\;\xi\int_{0}^{x}du\;\mathcal{K}(u;\nu,\epsilon)\,\bigl[\sigma_{z},\bigl[\sigma_{z},\,\langle\rho(x-u)\rangle\bigr]\bigr]\;,$ (17)
@@
$\xi\;\equiv\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\int_{0}^{\infty}du\;u^{-\nu}e^{-\epsilon/u}\;=\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\epsilon^{1-\nu}\,\Gamma(\nu-1)\;,$ (18)
and the unit‑normalised kernel is
@@
Equation (17) is the exact non‑Markovian master equation for neutrino propagation in red turbulence, valid non‑perturbatively for classical Gaussian noise.
@@
$\frac{d\rho_{12}}{dx}$ $=-i\bigl[\tilde{H}_{0},\rho\bigr]_{12}-\xi\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,\bigl[\sigma_{z},[\sigma_{z},\rho(x-u)]\bigr]_{12}$
$\simeq-i\Bigl(-\frac{1}{2}\cdot 2\rho_{12}\Bigr)-\xi\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,4\rho_{12}(x-u)$
$=i\rho_{12}(x)-4\xi\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,\rho_{12}(x-u).$ (25)
@@
Defining $\chi_{\nu}\equiv 2\xi$ and taking the Laplace transform,
@@
$s\widetilde{\rho}_{12}(s)-\rho_{12}(0)=i\widetilde{\rho}_{12}(s)-2\chi_{\nu}\,\tilde{\mathcal{K}}(s;\nu,\epsilon)\,\widetilde{\rho}_{12}(s),$ (27)
@@
$\bigl[s-i+2\chi_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]\,\widetilde{\rho}_{12}(s)\;=\;\rho_{12}(0).$ (28)
For notational clarity, we define the effective damping coefficient
$\kappa_{\nu}\equiv 4\xi,$ (29)
which absorbs the numerical factors arising from the double commutator structure. The off‑diagonal equation then takes the compact form
@@
Here $\kappa_{\nu}=4\xi$ represents the effective turbulence strength for decoherence, with $\xi$ given by Eq. (18).
The survival probability for an electron neutrino created at $x=0$ is
$P_{ee}(x)=\frac{1}{2}+\mathrm{Re}\!\bigl[\rho_{12}(x)\,e^{-ix}\bigr],$ (31)
where the phase $e^{-ix}$ originates from the matter‑basis oscillation factor $\exp(-i\Delta_{m}t)=\exp(-ix)$ in the high‑density limit. Taking the Laplace transform and using (30),
$\tilde{P}_{ee}(s)$ $=\frac{1}{2}\int_{0}^{\infty}e^{-sx}\,dx$
$+\frac{1}{2}\int_{0}^{\infty}e^{-sx}\,\bigl[\rho_{12}(x)e^{-ix}+\rho_{12}^{*}(x)e^{+ix}\bigr]dx$
$=\frac{1}{2}\cdot\frac{1}{s}+\frac{1}{2}\Bigl[\frac{\rho_{12}(0)}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s)-i}+\frac{\rho_{12}^{*}(0)}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s)+i}\Bigr].$ (32)
For an initial electron neutrino, $\rho_{12}(0)=\frac{1}{2}\sin 2\theta_{m}\simeq 0$ in the high‑density limit, but retaining the formal expression and setting $\rho_{12}(0)=\rho_{12}^{*}(0)=1/2$ (which amounts to an overall normalization) gives the compact result
$\tilde{P}_{ee}(s)\;=\;\frac{1}{2}\Bigl[\frac{1}{s}+\frac{1}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)-i}+\frac{1}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)+i}\Bigr]\;.\;$} (33)
Combining the three terms yields the all‑orders Laplace‑space expression quoted in equation (34).
Combining the three terms yields the compact, all‑orders expression
$\boxed{\;\tilde{P}_{ee}(s)\;=\;\frac{s\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]+1}{s\bigl\{\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]^{2}+1\bigr\}}\;.\;}$ (34)
This result, which emerges naturally from the non‑perturbative NZ treatment, encapsulates the complete interplay between coherent oscillations (poles at $s+\kappa_{\nu}\tilde{\mathcal{K}}(s)=\pm i$) and the algebraic decoherence induced by the long‑range correlations of red turbulence.

### v3

@@ Abstract
We develop an exact framework for neutrino decoherence in power-law correlated turbulent matter, as encountered in core-collapse supernovae. Employing the Nakajima–Zwanzig projection technique, we derive an exact non-Markovian master equation for the neutrino density matrix. For kernels $K(t)\propto t^{-\nu}$, the red-noise sector in our convention corresponds to $\nu<0$, while $\nu=1$ is the white-noise boundary. To treat ultraviolet singularities for $\nu\geq 1$ without spoiling the fractional structure, we use a renormalization prescription based on Hadamard finite parts and analytic continuation.
The exact Laplace-space solution for the survival probability is obtained. In the high-density matter basis relevant to supernovae, the solution is expressed through Mittag-Leffler functions, establishing a direct link to anomalous diffusion phenomena. For red spectra ($\nu<0$), the memory integral corresponds to a higher-order fractional operator.
Our work clarifies how spectral index, renormalization scale, and decoherence efficiency interrelate, providing a complete analytical description and practical tools for supernova neutrino simulations. The fractional calculus formulation reveals fundamental mathematical connections between neutrino flavor evolution and other systems governed by long-range temporal correlations.
[…]
@@ I Introduction
Core-collapse supernovae (CCSNe) represent one of the most energetic and complex astrophysical phenomena in the Universe, marking the cataclysmic death of massive stars ($M\gtrsim 8M_{\odot}$). These explosions not only seed the interstellar medium with heavy elements essential for planetary formation and life, but also give birth to neutron stars and black holes Janka (2012); Burrows et al. (2020); Müller et al. (2019). The detection of gravitational waves from binary neutron star mergers Abbott et al. (2017a, b) and the anticipated observation of neutrinos from the next Galactic supernova by next-generation detectors such as Hyper-Kamiokande Hyper-Kamiokande Proto-Collaboration et al. (2018) and DUNE Abi et al. (2021) underscore the importance of understanding these events through multi-messenger astronomy.
@@
Despite decades of theoretical and computational advancements, the precise explosion mechanism of CCSNe remains an open question Janka (2017); Burrows (2013); Foglizzo et al. (2015). Current three-dimensional simulations suggest that the neutrino-heating mechanism, aided by hydrodynamic instabilities such as the standing accretion shock instability (SASI) and neutrino-driven convection, can successfully revive the stalled shock in many progenitor models Abdikamalov et al. (2015); Couch and Ott (2013). Turbulence plays a crucial role in this process by enhancing the efficiency of neutrino heating and facilitating shock revival Couch et al. (2015); Calvert et al. (2025). Recent high-resolution simulations reveal that turbulent energy cascades follow Kolmogorov-like scaling in the inertial range Kolmogorov (1941), but the precise characterization of turbulence in supernova environments remains challenging Müller (2019b).
@@
Neutrinos, carrying away $\sim 99\%$ of the gravitational binding energy ($\sim 3\times 10^{53}$ erg), are the primary cooling agents and potential drivers of the explosion through the neutrino-heating mechanism Janka (2012); Müller (2019a). Their flavor evolution is profoundly affected by both coherent forward scattering (the Mikheyev-Smirnov-Wolfenstein (MSW) effect) Mikheyev and Smirnov (1985); Wolfenstein (1978) and collective oscillations due to neutrino-neutrino interactions Duan et al. (2010b, a); Galais et al. (2012). Recent studies have identified fast flavor instabilities that can occur on centimeter scales and nanosecond timescales Bhattacharyya and Dasgupta (2021); Dasgupta (2022); Ehring et al. (2023a); Fiorillo and Raffelt (2024), potentially altering the neutrino energy deposition and nucleosynthesis yields Nagakura (2023); Ehring et al. (2023b); Wang and Burrows (2025). Moreover, muon creation in supernova matter has been shown to facilitate neutrino-driven explosions Bollig et al. (2017), adding further complexity to the flavor evolution landscape.
@@
In addition to vacuum and matter effects, turbulence in the supernova envelope introduces stochastic density fluctuations that can significantly impact neutrino flavor evolution. The pioneering work of Loreti and Balantekin Loreti and Balantekin (1994) established the framework for treating neutrino oscillations in noisy media, demonstrating that density fluctuations can induce decoherence and modify survival probabilities. Subsequent studies have extended this analysis to various turbulent environments Kneller and Volpe (2010b); Reid et al. (2011); Kneller and Mauney (2013); Patton et al. (2014). For power-law correlated turbulence, the master equation acquires a non-Markovian character with memory kernels scaling as $K(t)\propto t^{-\nu}$, where $\nu$ is the spectral index determined by the turbulence power spectrum Kneller and Volpe (2010a); Borriello et al. (2014); Abbar (2021). Recently, Mukhopadhyay and Sen Mukhopadhyay and Sen (2024) investigated how turbulence signatures in supernova neutrinos could be probed in upcoming detectors, highlighting the observational relevance of this phenomenon.
@@
The mathematical treatment of neutrino decoherence in power-law correlated turbulent media presents significant challenges due to the non-Markovian nature of the evolution equations. The master equation with memory kernel $K(t)\propto t^{-\nu}$ naturally connects to fractional calculus, where the time convolution integral represents a fractional integral operator of order $1-\nu$ Jin (2021). For $\nu<1$, this corresponds to the Riemann-Liouville fractional integral, while for $\nu\geq 1$ one must isolate ultraviolet local pieces and absorb them through renormalization. Fractional calculus has found widespread application in describing anomalous diffusion, viscoelastic materials, and non-Markovian processes in open quantum systems Breuer and Petruccione (2007). In the context of neutrino physics, the emergence of Mittag-Leffler functions $E_{\alpha,\beta}(z)$ as natural solutions to fractional differential equations provides a powerful analytical framework Desai et al. (2016).
@@
While approximate and numerical treatments of neutrino propagation in turbulent media exist Loreti and Balantekin (1994); Kneller and Volpe (2010a)—ranging from white-noise models to power-law correlated fluctuations (numerical)—a complete analytical solution for the survival probability with generic power-law memory kernels has remained elusive. Previous approaches often relied on perturbative expansions or specialized limits, leaving the full mathematical structure unexplored. The power-law form of the memory kernel, $K(t)\propto t^{-\nu}$, suggests a natural connection to fractional calculus, yet this connection has not been fully exploited to obtain closed-form analytical solutions.
@@
Several critical gaps in the current understanding warrant systematic investigation: (i) the explicit formulation of the neutrino propagation problem within a fractional calculus framework; (ii) the renormalization required for spectral exponents $\nu\geq 1$ to handle ultraviolet divergences at short time scales ($t\to 0$); (iii) the derivation of exact analytical solutions for arbitrary $\nu$, expressed through special functions like Mittag-Leffler functions; (iv) the distinctive mathematical features arising for negative spectral exponents ($\nu<0$), which define red spectra; (v) the role of ultraviolet local terms in frequency renormalization; and (vi) the establishment of explicit connections between neutrino decoherence and anomalous diffusion phenomena through their shared fractional calculus structure.
@@
This work addresses these gaps by developing a comprehensive exact theory based on fractional calculus. We establish the precise correspondence between the neutrino master equation with power-law memory and fractional differential equations. Special attention is paid to the ultraviolet divergence of the correlation function for $\nu\geq 1$: we treat it by a renormalization prescription that absorbs the local divergent part into a frequency shift, while retaining the nonlocal power-law kernel. This yields exact, closed-form solutions valid for all $\nu$.
@@
The exact solution reveals a rich mathematical structure: for $\nu<1$, the memory integral corresponds to a Riemann-Liouville fractional integral of order $1-\nu$, while for red spectra ($\nu<0$) it becomes a higher-order fractional operator. The survival probability is expressed exactly in terms of Mittag-Leffler functions, which encode non-exponential relaxation with long memory. This formulation provides not only a complete analytical description but also efficient computational methods through established properties of fractional calculus and special functions.
@@
Our work demonstrates that the fractional calculus approach offers a unifying mathematical framework that connects neutrino decoherence in turbulent matter with broader phenomena in statistical physics, including anomalous diffusion and viscoelastic relaxation. This perspective yields new insights into the interplay between spectral index, renormalization scale, and decoherence efficiency, providing both fundamental understanding and practical tools for supernova neutrino simulations.
@@
In this work, we establish the complete mathematical formalism for the exact theory of neutrino decoherence in turbulence with red spectra (defined by $\nu<0$). We provide a rigorous derivation of the master equation for the neutrino density matrix via an open quantum systems approach, obtaining the exact non-Markovian equation through the Nakajima–Zwanzig projection technique Breuer and Petruccione (2007). Furthermore, we derive the exact solution for the electron neutrino survival probability in Laplace space. For the red-spectral sector, we develop a fractional calculus framework, yielding a fractional differential equation exactly solvable in terms of Mittag-Leffler functions. This fractional formulation not only provides efficient computational methods but also uncovers profound mathematical links between neutrino flavor evolution and other physical systems characterized by power-law correlations and anomalous dynamics.
@@
This paper is organized as follows: Section II reviews the exact Laplace-space solution using an auxiliary short-time regulated representation of the kernel. Section III develops the renormalized analytic solution via fractional calculus. Section IV summarizes our findings and outlines directions for future research. In a companion paper Bao et al. (2026), we extend this analysis to include collective neutrino oscillations in turbulent environments.
@@ II.1 Physical Picture and Assumptions
We begin by outlining the physical scenario and the key assumptions underlying our derivation. The propagation of neutrinos through turbulent matter in core-collapse supernovae involves multiple scales:(a)Microscopic scale ($\sim 10^{-12}$ cm): Neutrino flavor oscillations governed by quantum mechanics. (b)Turbulence inertial range ($10^{2}$$10^{6}$ cm): Compressible hydrodynamic turbulence with power-law scaling Frisch (1995). and (c) Supernova scale ($\sim 10^{8}$ cm): Overall system size.
@@ II.2.1 Obukhov-Corrsin Spectrum and Dimensionless Formulation
Let us begin by introducing dimensionless variables for a formal simplification. Let $n_{0}$ be a reference electron number density (typically the mean density $\bar{n}_{e}$), and $T_{0}$ a reference time (to be chosen as $1/\Delta_{m}$, where $\Delta_{m}=\frac{G_{F}n_{0}}{\sqrt{2}}\cos 2\theta_{m}$ is the matter-level splitting of neutrinos). In natural units ($\hbar=c=1$), length, time, and energy satisfy $L]=[T]=[E]^{-1}$. We define dimensionless coordinates and density fluctuations:
@@
which is itself dimensionless. For fully developed power-law turbulence, the correlation function exhibits power-law scaling in the inertial range. While large-scale correlations are naturally limited by the system size, the small-scale behavior requires careful treatment due to potential divergences at $\tau\to 0$. In this section we use an auxiliary short-time regulated representation of the kernel; the renormalized formulation is developed in Sec. III.
II.2.2 Auxiliary Short-Time Regulated Correlation Function
The power-law correlation $K(\tau)\propto\tau^{-\nu}$ diverges as $\tau\to 0$ for $\nu\geq 1$, representing an ultraviolet singularity. For the intermediate derivation in this section, we introduce an auxiliary short-time regulator at the dissipation scale $\tau_{\text{diss}}=l_{\text{diss}}/c$, where $l_{\text{diss}}$ is the Kolmogorov dissipation scale:
@@ II.2.2 Auxiliary Short-Time Regulated Correlation Function
For $\nu<1$, the correlation function is integrable at $\tau=0$ even without this regulator; this includes the red-noise sector $\nu<0$ adopted in this work. The normalization constant $A$ is chosen such that $\int_{0}^{\infty}\tilde{\mathcal{C}}(\tilde{\tau})d\tilde{\tau}=1$ when considering the effective damping rate.
@@ II.2.3 Exact Non-Markovian Master Equation from Nakajima–Zwanzig Projection
We now present a rigorous, non-perturbative derivation of the master equation governing neutrino decoherence in turbulent matter (with red spectra corresponding to $\nu<0$) using the Nakajima–Zwanzig (NZ) projection operator formalism Breuer and Petruccione (2007). The essential advantage of the NZ approach is that it treats the memory kernel exactly for classical Gaussian noise, automatically resumming all cumulants that contribute to the phase‑diffusion process.
@@
$-\;\xi_{\nu}\int_{0}^{x}du\;\mathcal{K}(u;\nu,\epsilon)\,\bigl[\sigma_{z},\bigl[\sigma_{z},\,\langle\rho(x-u)\rangle\bigr]\bigr]\;,$ (17)
@@
$\xi_{\nu}\;\equiv\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\int_{0}^{\infty}du\;u^{-\nu}e^{-\epsilon/u}\;=\;\frac{G_{F}^{2}\,n_{0}^{2}\,A}{2\Delta_{m}^{2}}\,\epsilon^{1-\nu}\,\Gamma(\nu-1)\;,$ (18)
and
[…]
where $\kappa_{\nu}$ characterizes the turbulence-induced decoherence rate and combines $G_{F}^{2}n_{0}^{2}$, the matter-oscillation scale $\Delta_{m}^{-2}$, and the spectral-index dependence through the correlation integral. The unit‑normalised kernel is
@@
Equation (17) is the exact non‑Markovian master equation for neutrino propagation in turbulence with red spectra ($\nu<0$), valid non‑perturbatively for classical Gaussian noise.
@@
$\frac{d\rho_{12}}{dx}$ $=-i\bigl[\tilde{H}_{0},\rho\bigr]_{12}-\xi_{\nu}\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,\bigl[\sigma_{z},[\sigma_{z},\rho(x-u)]\bigr]_{12}$
$\simeq-i\Bigl(-\frac{1}{2}\cdot 2\rho_{12}\Bigr)-\xi_{\nu}\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,4\rho_{12}(x-u)$
$=i\rho_{12}(x)-4\xi_{\nu}\!\int_{0}^{x}\!du\;\mathcal{K}(u;\nu,\epsilon)\,\rho_{12}(x-u).$ (26)
@@
Using Eq. (19), namely $4\xi_{\nu}=\kappa_{\nu}$ (equivalently $2\xi_{\nu}=\kappa_{\nu}/2$), and taking the Laplace transform,
@@
$s\widetilde{\rho}_{12}(s)-\rho_{12}(0)=i\widetilde{\rho}_{12}(s)-\kappa_{\nu}\,\tilde{\mathcal{K}}(s;\nu,\epsilon)\,\widetilde{\rho}_{12}(s),$ (28)
@@
$\bigl[s-i+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]\,\widetilde{\rho}_{12}(s)\;=\;\rho_{12}(0).$ (29)
The off‑diagonal equation then takes the compact form
@@
Here $\kappa_{\nu}$ is the unique turbulence-strength parameter used in the following equations, and it is defined by Eq. (19).
The survival probability for an electron neutrino is obtained by projecting the matter-basis density matrix onto the flavor state:
$P_{ee}(x)=\cos^{2}\theta_{m}\,\rho_{11}(x)+\sin^{2}\theta_{m}\,\rho_{22}(x)+\sin 2\theta_{m}\,\mathrm{Re}\!\bigl[\rho_{12}(x)\bigr].$ (31)
For an initial electron neutrino at $x=0$,
$\rho_{11}(0)=\cos^{2}\theta_{m},\qquad\rho_{22}(0)=\sin^{2}\theta_{m},\qquad\rho_{12}(0)=\frac{1}{2}\sin 2\theta_{m}.$ (32)
In the high-density limit used here, the diagonal populations decouple and remain constant, so
$P_{ee}(x)=1-\frac{1}{2}\sin^{2}2\theta_{m}+\sin 2\theta_{m}\,\mathrm{Re}\!\bigl[\rho_{12}(x)\bigr].$ (33)
Taking the Laplace transform and using (30),
$\tilde{P}_{ee}(s)$ $=\frac{1-\frac{1}{2}\sin^{2}2\theta_{m}}{s}+\frac{\sin^{2}2\theta_{m}}{4}\Biggl[\frac{1}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)-i}+\frac{1}{s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)+i}\Biggr].$ (34)
Equivalently, the compact all-orders form is
$\boxed{\;\tilde{P}_{ee}(s)=\frac{1-\frac{1}{2}\sin^{2}2\theta_{m}}{s}+\frac{\frac{1}{2}\sin^{2}2\theta_{m}\,\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]}{\bigl[s+\kappa_{\nu}\tilde{\mathcal{K}}(s;\nu,\epsilon)\bigr]^{2}+1}\;.\;}$ (35)
which enforces $P_{ee}(0)=1$ by construction (initial-value theorem) and reduces to the standard coherent matter-oscillation limit as $\kappa_{\nu}\to 0$. This result, which emerges naturally from the non‑perturbative NZ treatment, encapsulates the complete interplay between coherent oscillations (poles at $s+\kappa_{\nu}\tilde{\mathcal{K}}(s)=\pm i$) and the algebraic decoherence induced by long-range correlations in the red-spectral sector ($\nu<0$).